# NoiseBench: Benchmarking the Impact of Real Label Noise on Named Entity Recognition

### v1

@@ 4.2 Results
Clean subset has high potential. The upper bound of training only on clean examples is the best achievable result for all types of noise. This was to be expected, since in this scenario the labels are not contaminated by any noise. This indicates that noise-robust approaches focusing on identifying the clean subset have high potential. Oracle stopping, on the other hand, does not achieve the same level of performance as the oracle subset, while only slightly outperforming the FLERT baseline. This is in line with our findings in Experiment 2 that the early-learning generalization phase is skipped when training with real noise, indicating that early stopping approaches have less potential.
Small benefit of noise-robust approaches. Evidently, there is no single best approach for all noise types. While for all noise types some of the noise-robust approaches are able to outperform the baseline, they do so only slightly, with the exception of BOND, on the Distant and Crowd variants. Still, their performance is far below the upper bound. This raises the issue of trade-offs of existing noise-robust learning approaches, since they often require additional hyperparameter tuning or incur computational costs, but only lead to slight improvements over the baseline in the presence of real noise.
@@ 5 Related work
There are a few benchmarks for learning with label noise and related topics. The WRENCH benchmark Zhang et al. (2021b) focuses only on weak supervision labels for multiple tasks, with the emphasis on combining multiple weak labelling sources. Klie et al. (2023) compare a large number of methods for the detection of annotation errors. Multiple tasks are included, including NER on CoNLL-03, where they evaluate the detection of expert errors, concluding that most approaches are not successful at this. Similarly, Chong et al. (2022) evaluate annotation error detection on datasets with noise only from crowdsourced labels, for part-of-speech tagging and natural language inference tasks. Liu et al. (2022) propose a benchmark for text classification under label noise, where they re-annotate an existing sentiment classification dataset and construct noisy label sets according to annotator disagreements; however, they do not publish these label sets. NoisyWikiHow, a benchmark for intention identification has also been presented Wu et al. (2023), where the authors propose a method to simulate realistic noise that closely imitates human errors by producing heterogeneous and instance-dependent errors. For NER in Estonian, Hedderich et al. (2021) introduce the NoisyNER, which includes multiple noise levels obtained from distant supervision approaches with varying quality.
@@ 6 Conclusion
In this paper, we address the issue of label noise in the NER task. We introduce a new benchmark, based on the commonly used NER dataset CoNLL-03 in English, for evaluating the impact of 6 distinct types of real label noise on the same set of sentences, with varying degrees of difficulty.
We demonstrate that real noise causes transformer-based language models to immediately memorize the noise pattern, making real label noise a more challenging problem than simulated label noise, even in the case of oracle class-dependent noise informed by the characteristics of real noise.
We further investigate this by evaluating popular approaches aimed at mitigating this issue and showcasing their limitations. Our experiments indicate that current methods fall far short of what can be potentially achieved on almost all of the noisy datasets in NoiseBench, leaving room for improvement. We hope that NoiseBench aids other researchers in the further development of more effective noise-robust approaches and in characterizing the differences between types of real noise.
@@
We identify two limitations in our study.
First, it focuses on the scenario when the entire available dataset could be noisy and we do not have access to a small, high-quality labelled, data subset. While this is a certainly scenario which reflects a large number of real-world cases, it could be argued that in some situations it is realistic to have the resources to ensure a subset of the data is clean, with high-quality annotations. However, when this is the case, Zhu et al. (2023) showcased that this clean data would be better utilized by directly fine-tuning the models on it, instead of using it for validation. Therefore, we argue that this alternative setup is not particularly useful for the evaluation of label-noise-robust approaches.
Second, this benchmark is limited to the English language, because it is dependent on the availability of various labelling sources for the widely used CoNLL dataset. Learning under label noise is a greater challenge in domains where data is limited, which is why investigating this problem for low-resource languages would be beneficial for the research community. Nevertheless, we believe the findings regarding noise robustness obtained with our benchmark would be transferable to data in other languages.
@@ Acknowledgements
Elena Merdjanovska and Alan Akbik are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135. Ansar Aynetdinov and Alan Akbik are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Emmy Noether grant “Eidetic Representations of Natural Language” (project number 448414230).
Appendix A Baseline for uniform noise
@@
In Table 4 the results from Experiment 1 for uniform noise are shown. We can see the the model is quite robust to uniform noise and that it results in higher test performance, when compared to real noise of the same level. The average difference in F1 scores in 17 percentage points, which is why we focus on the more realistic oracle class-dependent noise simulation method in the remainder of the results of Experiments 1 and 2.
@@
Appendix B Generalization to unseen entities
(a) Real noise
(b) Simulated class-dependent noise
Figure 3: Comparison of F1 scores on different subsets of entities in the test set: all, seen (clean), seen (noisy) and unseen.
We evaluate the generalization of the baseline model from Experiment 1 (Section 3.3) by calculating F1 scores for seen and unseen entities separately, and further distinguishing seen entities according to whether their label in the training set was clean or noise-corrupted. On Figure 3(a) we observe that for the Expert and Crowd++ noise types, the performance for the seen (clean) entities and the unseen entities is comparable, which indicates the model has the ability to generalize to unseen entities well. The overall drop in performance when training with these noisy training splits is largely due to the low performance on the seen (noisy) entities. However, this detrimental effect of memorization is limited to the seen errors and does not affect the performance on unseen entities much, for noise levels up to 15%. As for the remaining training splits with noise levels of 30% and above, we can see that noise affects the performance on unseen entities as well. Furthermore, on Figure 3(b), we see that for simulated noise, for Expert and Crowd++ , the performance on seen noisy entities is close to the performance on seen clean entities. This means that at low noise levels, the models are able to disregard the noisy patterns and predict the same entities correctly when they appear in the test set, instead of memorizing the noisy label.
Appendix C Memorization of Crowd, Weak and LLM training variants
The memorization plots from Experiment 2, as an extension of Section 3.4 for the Crowd, Weak and LLM dataset variants are shown in Table 4
@@ Appendix C Memorization of Crowd, Weak and LLM training variants
Figure 4: Comparison of model performance during extended training. The top row shows models fine-tuned on label sets with real noise, while the bottom row models fine-tuned on a corresponding simulated noisy label set (class-dependent noise). The graphs show both the F1 score on the noisy training labels and on the clean training labels, for 3 different noise types. The plots are averages of 3 runs.
. We again observe immediate memorization of real noise and delayed memorization of simulated noise.
Appendix D Implementation details
D.1 Confident learning
D.2 Co-regularization
We use two classifiers and perform a hyperparameter sweep as suggested by the authors.
D.3 BOND
Since we cannot perform early stopping on a clean validation set, which is suggested for both stages by the authors, we rely on the findings reported by Tänzer et al. (2022) and stop the first stage after the first 4 epochs. The second stage is limited to 6 epochs in order to reproduce the same training length as in the FLERT baseline. We adopt the hyperparameters reported as best by the authors for the CoNLL-03 dataset, and use hard pseudo-labels in the second stage.

### v2

@@ 4.2 Results
Identifying a clean subset has highest potential. The upper bound of training only the clean subset of each noisy split (see "Oracle subset" in Table 3) achieves the best scores of all upper bounds. This makes intuitive sense as training is performed only over fully clean sentences, albeit a smaller subset of the full training data as all noisy sentences are filtered out. Similarly, we find strong improvements for the "Additional clean data" upper bound. Oracle stopping, on the other hand, does not achieve the same level of performance as the oracle subset, only slightly outperforming the FLERT baseline. This is in line with our findings in Experiment 2 that the early-learning generalization phase is skipped when training with real noise. This indicates that noise-robust learning approaches that target early stopping have little potential.
Small benefit of noise-robust approaches. Evidently, there is no single best approach for all noise types. For each noise type, at least one noise-robust approach outperforms the baseline, however on average most of them are comparable to it. Only MSR outperforms the baseline averaged over all noise types, bringing improvements for Crowd++, Distant, Weak and LLM. Additionally, L2R works well for LLM noise and BOND for Crowd. Still, the performance is far below the upper bound. This raises the issue of trade-offs of existing noise-robust learning approaches, since they often require additional hyperparameter tuning or incur computational costs, but only lead to slight improvements in the presence of real noise.
5 Ablation: NoiseBench for German
Using the German sentences in CoNLL-03, we created a noisy label benchmark for German to confirm our findings for a different language. Following the English counterpart described in Section 2, it consists of (1) a noise-free test split to evaluate trained models and (2) three variants of the training split, where two are annotated with different types of noise and one is the ground truth. The two types of noise include Expert labels, with 16.2% noise and LLM labels, with 54% noise. More details can be found in Appendix A.1.
5.1 Experimental Results
Clean Expert LLM Avg.
Baseline 90.24±0.2 79.02±0.4 57.86±0.4 75.7
Upper bounds
Oracle subset - 83.11±0.6 82.72±0.6 82.9
Oracle stopping 90.50±0.2 79.48±0.3 61.81±0.9 77.3
Additional clean data 89.86±0.9 82.85±1.7 69.50±1.4 80.7
Noise-robust learning
Confident learning 90.00±0.3 79.57±0.3 58.03±0.2 75.9
CrossWeigh 90.11±0.3 78.32±0.3 57.50±0.8 75.3
L2R 81.45±3.1 74.14±0.8 53.07±1.6 69.6
Co-regularization 88.50±0.2 78.49±0.2 54.47±0.5 73.8
BOND 86.53±0.3 77.56±0.5 55.89±0.6 73.3
MSR 85.34±0.5 76.42±0.4 64.00±0.7 75.3
Table 4: German variant: Performance of noise-robust approaches on the Clean test set, when training on each training split variant. Results are expressed in terms of F1 score. Each score is averaged over 3 runs.
Table 4 shows the results of the noise-robust approaches and upper bounds when training on the German datasets. More experimental details and results can be found in Appendix A.2.
Oracle subset score reaches an upper limit. Regarding the upper bounds, we see that the performance of the oracle subset of Expert and LLM is close, meaning that the 4000555See Appendix C for the size of the oracle subset. clean sentences in the LLM subset are already enough to reach an F1 score over 82. Despite having more samples, the Expert subset does not result in a much higher score. This could signify that the remaining sentences, not included in the Expert subset, are difficult examples necessary to properly learn the task.
Poor performance of noise-robust approaches. Regarding the noise-robust approaches, only Confident Learning is able to match and slightly outperform the baseline. All other methods mainly perform poorly on the German dataset, even below the baseline, with the exception of the improvement brought by MSR on the LLM dataset.
@@ 6 Related Work
There are a few benchmarks for learning with label noise and related areas. The WRENCH benchmark Zhang et al. (2021b) focuses only on weak supervision labels for multiple tasks, with the emphasis on combining multiple weak labelling sources. Klie et al. (2023) compare a large number of methods for the detection of annotation errors. Multiple tasks are included, including NER on CoNLL-03, where they evaluate the detection of expert errors, concluding that most approaches are not successful at this. Similarly, Chong et al. (2022) evaluate annotation error detection on datasets with noise only from crowdsourced labels, for part-of-speech tagging and natural language inference tasks. Liu et al. (2022) propose a benchmark for text classification under label noise, where they re-annotate an existing sentiment classification dataset and construct noisy label sets according to annotator disagreements; however, they do not publish these label sets. NoisyWikiHow, a benchmark for intention identification has also been presented Wu et al. (2023), where the authors propose a method to simulate realistic noise that imitates human errors by producing heterogeneous and instance-dependent errors. For NER in Estonian, Hedderich et al. (2021) introduce the NoisyNER, which includes multiple noise levels obtained from distant supervision approaches with varying quality. MultiCoNERv2 Fetahu et al. (2023) addresses textual noise in the input data itself (e.g. typos), instead of label noise.
@@ 7 Conclusion
In this paper, we address the issue of label noise in the NER task. We introduce a new benchmark, based on the commonly used NER dataset CoNLL-03, for evaluating the impact of 6 distinct types of real label noise on the same set of sentences, with varying degrees of difficulty.
We demonstrated that real noise causes transformer-based language models to immediately memorize the noise pattern, making real label noise a more challenging problem than simulated label noise, even in the case of oracle class-dependent noise informed by the characteristics of real noise.
We further presented an evaluation of popular noise-robust learning approaches. Our experiments indicate that current methods fall far short of what can potentially be achieved on the noise types in NoiseBench and that approaches that focus on automatically identifying a clean subset of labels have the highest potential. We hope that NoiseBench aids other researchers in the further development of more effective noise-robust approaches.
@@
This paper focuses on the scenario when the entire available dataset could be noisy and we do not have access to a small, high-quality labelled, data subset. While this is a certainly scenario which reflects a large number of real-world cases, it could be argued that in some situations it is realistic to have the resources to ensure a subset of the data is clean, with high-quality annotations. However, when this is the case, Zhu et al. (2023b) showcased that this clean data would be better utilized by directly fine-tuning the models on it, instead of using it for validation. Therefore, we argue that this alternative setup is not particularly useful for the evaluation of label-noise-robust approaches.
@@ Acknowledgements
We thank all reviewers for their valuable comments. Elena Merdjanovska and Alan Akbik are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135. Ansar Aynetdinov and Alan Akbik are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Emmy Noether grant “Eidetic Representations of Natural Language” (project number 448414230).
Appendix A NoiseBench for German: Additional Details
Following are additional details about the German version of NoiseBench, as well as results from Experiment 2 for this dataset.
A.1 Overview
The training split contains 12,705 sentences from 553 documents, covering 10,008 entity mentions. The test split contains 3,160 sentences from 155 documents, covering 3,051 entity mentions.
A.1.1 Noise-Free Data
As noise-free labels for the German part of CoNLL-03, we take the updated annotations from 2006666More details about the revision of the labels can be found in the ner.tz file, downloaded from https://www.clips.uantwerpen.be/conll2003/ner/, more specifically in the /ner/etc.2006/revision.txt and ner/etc.2006/guide.pdf files.. This updated label set is considered ground-truth by the research community.
When compared to the original CoNLL-03 German labels, he most changes in this label set are for the MISC class, most notably the removal of adjectives derived from names. This can alternatively be considered an update in annotation guidelines. With this, it should be noted that for the German dataset we do not have access to labels with verified high quality, as we do for the English counterpart with the CleanCoNLL labels.
A.1.2 Expert Errors
Similar as for English, we take the original CoNLL-03 labels Tjong Kim Sang and De Meulder (2003) as labels with expert errors. This results in a noise share of 16.2%.
A.1.3 LLM Teacher Models
Similar as for English, we use GPT3.5 to create a noisy version of the training split annotated by an LLM. This results in a high noise share of 54%.
A.1.4 Statistics
When compared to the noise shares for Expert and LLM in NoiseBench in Table 1 (5.5% and 45.6%), the noise shares for German are higher (16.2% and 54% respectively). This is due to LLMs performing more poorly on languages other than English, as well as due to fewer research efforts focusing on re-annotating and cleaning the German part of CoNLL-03, resulting in less consistent labels.
We see that most of the errors in German are non-entity mentions. Overall, we also note that type errors and partial matches are much less prominent here (less than 10%), even though they formed a larger part of the errors in NoiseBench.
#Entities %Errors
Expert 16.2 98.2 83.8 11852 9156 14.5 73.0 6.1 6.4
LLM 54.0 92.1 46.0 16526 6102 19.8 69.9 3.2 7.0
Table 5: Overview of the noisy training splits in NoiseBench for German. The table shows the noise level, the micro-averaged token-level F1 score (F1token), micro-averaged entity-level F1 (F1entity), the number of entities (Total), number of correct entities (Correct) and share of each error type: missing mentions (Missing (FN)), non-entity mentions (Non-entity (FP)), wrong type (Type) and partial matches (Partial). All metrics are in comparison to the Clean split.
A.2 Experiments
A.2.1 Validation Split
We take the last 96 documents from the training split to serve as a validation set, corresponding to roughly 17% of all sentences.
A.2.2 Experiment 2
We performed Experiment 2 for the German version of NoiseBench, where the goal is to observe the memorization of label noise. The resulting graphs for both noise types are shown in Figure 4. We observe a similar behaviour as in the English part of NoiseBench, where the real noisy datasets are memorized immediately.
(a) Expert - 16.4% noise
(b) LLM - 54.7% noise
Figure 4: Comparison of model performance during extended training, for the German dataset.
Appendix B Implementation Details
In all our experiments with noise-robust methods we use the same xlm-roberta-large transformer as in the FLERT baseline with a batch size of 32, except for L2R, for which we used a batch size of 16 due to VRAM constraints.
B.1 Confident Learning
B.2 Co-Regularization
We perform a hyperparameter sweep as suggested by the authors, and choose the best performing ones on the validation set of the respective noise type.
B.3 BOND
In our experiments we found that limiting BOND’s first stage of training to 1 epoch is not enough for optimal performance, hence why we rely on the findings reported by Tänzer et al. (2022) and stop the first stage after the first 3 epochs. The second stage is limited to 7 epochs in order to reproduce the same training length as in the FLERT baseline. We update the teacher model in the second stage of training every 2 epochs as suggested by the authors for the CoNLL-03 dataset, and use hard pseudo-labels in the second stage, which we found to outperform soft pseudo-labels in our experiments.
B.4 CrossWeigh
We ran the CrossWeigh framework with 5 folds and 3 iterations, because according to the ablation experiments ran by the authors, higher numbers did not bring significant performance improvements. For a fair comparison, we adjust the CrossWeigh framework to use transformer fine-tuning as a base model. For the final training run using the sample weights, we used the same FLERT approach as in the baseline.
B.5 L2R
We rely on the implementation provided by Zhu et al. (2023b), test out two meta-learning rates while keeping the model learning rate fixed at 5e-6, and perform the validation step every 0.1 epoch with a patience of 10 validation steps.
B.6 MSR
We used the implementation provided by the authors Zhu et al. (2023a) and the hyperparameters they selected for CoNLL, as stated in their paper. For the German dataset, we used xlm-roberta-base as a multilingual model.
B.7 Upper Bound: Additional Clean Data
This upper bound assumes an additional small dataset with high-quality labels is available. We fix this number to 100 sentences, which are randomly chosen from the validation split (otherwise not used to train the models). This training setting first fine-tunes the baseline model for 10 epochs, and then continues fine-tuning only on the small clean dataset for 5 more epochs.
B.8 GPT3.5
To obtain an LLM-annotated variant of the training splits, for NoiseBench we used gpt-3.5-turbo-0613, while for German we used gpt-3.5-turbo-0125.
Appendix C Size of Oracle Subsets
In Table 6, the number of clean sentences in the oracle subset for each noisy variant is shown. These oracle subsets are used as performance upper bounds, as explained in Section 4.1.1.
% of all sentences Oracle subset size
NoiseBench split
Clean 100.0 4879
Expert 92.6 4483
Crowd++ 79.4 3786
Crowd 55.3 2554
Distant 59.7 2728
Weak 49.6 2294
LLM 38.2 1705
German split
Clean 100.0 10824
Expert 81.6 8827
LLM 37.8 4095
Table 6: Details about the oracle subset used as an upper performance bound. The table shows the percentage of clean sentences and the absolute number, for each noise type.
Appendix D Baseline for Uniform Noise
@@
Table 7 shows the results from Experiment 1 for uniform noise. We can see the the model is quite robust to uniform noise and that it results in higher test performance, when compared to real noise of the same level. The average difference in F1 scores in 17 percentage points, which is why we focus on the more realistic oracle class-dependent noise simulation method in the main results of Experiments 1 and 2.
@@
Appendix E Memorization of Crowd, Weak and LLM Noise
Figure 5 is an extension of Section 3.4 and shows the memorization plots from Experiment 2 for the Crowd, Weak and LLM dataset variants. We again observe immediate memorization of real noise and delayed memorization of simulated noise.
@@ Appendix E Memorization of Crowd, Weak and LLM Noise
Figure 5: Comparison of model performance during extended training, for Crowd, Weak and LLM from NoiseBench. The top row shows models fine-tuned on label sets with real noise, while the bottom row models fine-tuned on a corresponding simulated noisy label set. The graphs show both the F1 score on the noisy training labels and on the clean training labels, for 3 different noise types. The plots are averages of 3 runs.
Appendix F Additional Experiments on Memorization
In addition to the main Experiment 2, we ran two ablation experiments regarding memorization.
F.1 Effect of Pre-training on Memorization
The first ablation compares fine-tuning a pretrained model and a model with randomly initialized weights. Figure 6 shows this comparison during an extended training run for the Crowd++ training variant, where we used DistilBERT (learning rate of 5e-05). We can see that even without pretraining, the model starts overfitting to the noisy labels and we can observe a large gap between the performance on the clean and noisy labels.
(a) Pre-trained model
(b) Random initialization
Figure 6: Memorization of label noise in DistilBert, using the pretrained model and a model with randomly initialized weights. The experiment was run for one noise type - Crowd++.
F.2 Memorization in a Smaller Model
The second experiment investigates memorization when fine-tuning a smaller model, DistilRoBERTa, because the model we use in the main experiments, XLM-RoBERTa-Large, is quite large. Figure 7 shows an extended training run for three noise types in NoiseBench. We observe the same patterns of immediate and delayed memorization as with the larger model.
Expert Crowd++ Distant
Real
(a) Real Expert  - 5.5% noise
(b) Real Crowd++  - 18% noise
(c) Real Distant  - 31.3% noise
Simulated
(d) Simul. Expert  - 6% noise
(e) Simul. Crowd++  - 15% noise
(f) Simul. Distant  - 39% noise
Figure 7: Comparison of model performance during extended training with a smaller model, Distil-RoBERTa. The top row shows models fine-tuned on label sets with real noise, while the bottom row models fine-tuned on a corresponding simulated noisy label set. The graphs show the F1 score on the noisy training labels and on the clean training labels, for 3 noise types. The plots are averages of 3 runs.
Appendix G Extended Performance Metrics
In this section we provide extended metrics of the predictive performance of the baseline FLERT method. These metrics and analysis correspond to Experiment 1 from Section 3.3.
G.1 Analysis of Test Errors
We can characterize the model predictions in a similar way as we characterized the different types of errors in Table 1 and Figure 1. Table 8 shows how representative different types of prediction errors are, expressed as a percentage of all errors. We can see that with Expert noise, a majority of the mistakes are wrong entity types. Furthermore, for the Crowd and Distant dataset versions, the largest number of errors is due to missing entities, while for the Weak and LLM datasets, the errors are mostly non-entities or wrong type. This is in line with the characteristics of the noisy datasets themselves, described in Table 1.
For German we make similar observations. For the clean variant, most errors are missing entities. However, for the two noisy variants, which include a large number of noisy non-entity annotations as seen in Table 5, the majority of prediction errors are also non-entity mentions.
% Prediction errors
Missing Non-entity Type Partial
NoiseBench split
Clean 13.9 25.4 29.6 31.1
Expert 12.2 14.4 54.1 19.3
Crowd++ 27.0 11.9 42.1 18.9
Crowd 55.2 6.4 25.3 13.1
Distant 64.6 8.2 14.4 12.8
Weak 14.2 24.5 49.5 11.8
LLM 20.5 34.9 38.3 6.4
German split
Clean 39.9 26.4 16.7 16.9
Expert 17.8 61.9 11.7 8.6
LLM 26.7 61.5 4.8 6.9
Table 8: Overview of the percentage of different types of prediction errors
We also examined the confusion matrices of the predictions. We were able to identify some patterns, regarding which types of errors are more prone to memorization. For most noise types in English NoiseBench, the largest number of prediction mistakes (out of the strings previously seen with a noisy label in the training set) were missing ORG and MISC entities, as well as ORG misclassified as LOC. These mistakes were present in a large number consistently across noise types.
However, we also observed a large number of missing ORG and MISC entities in the predictions when using the clean training set, which indicates that this is an inherently difficult pattern, even when noise is not present. On another hand, the pattern of misclassifying ORG as LOC does not happen when clean data is available. Therefore we can conclude that when this type of noisy pattern is present in the training set, the models are not able to recognize it as noise and are not robust to it.
G.2 Per-Class Metrics
We provide per-class metrics for a more extensive evaluation of the performance of the baseline method, for both German and English, in Table 9.
We can see that both precision and recall for the MISC class are generally lower than the other classes. This is especially noteworthy in the Expert label set of the German split, which does not have a high noise share, but it does have very low performance on MISC. This is however expected, as most of the noisy labels in this label set are related to MISC entities.
LOC ORG PER MISC
Prec. Recall #Ent. Prec. Recall #Ent. Prec. Recall #Ent. Prec. Recall #Ent.
NoiseBench split
Clean 93.9 ±0.6 93.5 ±0.4 1413 93.5 ±0.1 94.6 ±0.1 1909 99.0 ±0.2 99.1 ±0.1 1591 82.9 ±1.0 86.2 ±1.0 812
Expert 81.1 ±0.1 95.5 ±0.6 1413 92.0 ±0.6 81.9 ±0.3 1909 98.4 ±0.7 99.0 ±0.3 1591 85.3 ±0.9 81.4 ±1.1 812
Crowd++ 75.9 ±0.6 94.4 ±0.4 1413 91.8 ±0.3 75.4 ±0.8 1909 97.6 ±0.5 97.8 ±0.3 1591 86.5 ±0.5 70.8 ±0.1 812
Crowd 65.5 ±0.4 90.7 ±1.2 1413 83.7 ±0.8 44.5 ±0.3 1909 93.6 ±1.9 71.7 ±1.2 1591 84.4 ±0.4 47.5 ±0.6 812
Distant 83.8 ±0.7 74.8 ±0.2 1413 85.3 ±1.3 55.9 ±1.0 1909 75.7 ±1.7 84.8 ±0.8 1591 98.6 ±1.4 13.8 ±3.1 812
Weak 52.5 ±0.2 93.1 ±0.1 1413 49.9 ±0.2 34.6 ±0.3 1909 88.5 ±0.3 88.9 ±1.0 1591 84.2 ±0.9 57.3 ±0.9 812
LLM 52.7 ±0.3 84.6 ±0.5 1413 57.7 ±0.9 45.4 ±0.5 1909 95.4 ±0.6 98.3 ±0.1 1591 12.7 ±0.4 11.8 ±0.4 812
German split
Clean 93.0 ±0.7 90.8 ±0.2 1051 80.3 ±0.6 82.3 ±0.3 584 96.6 ±0.4 97.1 ±0.2 1210 77.1 ±1.7 56.5 ±0.6 206
Expert 88.4 ±0.6 83.3 ±0.9 1051 67.1 ±0.8 84.0 ±0.2 584 96.9 ±1.0 96.0 ±0.4 1210 13.1 ±0.3 40.6 ±0.8 206
LLM 63.9 ±0.4 70.3 ±0.9 1051 37.3 ±0.7 82.8 ±1.4 584 66.8 ±0.2 63.7 ±0.6 1210 10.4 ±0.6 18.4 ±1.6 206
Table 9: Per-class metrics of the predictions on the Clean test set